# Adversarial Agent Collaboration for C to Rust Translation

### v1

@@ 4.2 Evaluation Result on Micro Benchmark
Figure 2: Overall correctness (pass rate) achieved by ACToR on micro benchmark across different settings compared with naive baseline, on 3 different agent-model settings.
@@
Correctness & Adaptability. We first test ACToR on the micro benchmark with different agents and LLMs. To analyze the correctness improvement, we compare ACToR approach with a naive baseline on pass rate. The experiment results per program are shown in Figure 2. All translations from ACToR and the naive baseline are 100% safe and free of unsafe sections. ACToR consistently improves the translation correctness on all 3 agent-model settings. With Claude Code, the validation pass rate rises from an average of 79.3% (one-shot naive baseline) to 92.1% after 10 iterations. For Mini-SWE-Agent, pass rates increase from 81.9% to 90.7% with Sonnet-4 and from 84.1% to 86.8% with GPT-5 mini on average. Across agent-LLM settings, the closed-source Claude Code achieves the highest final pass rate. We therefore use Claude Code + Claude-Sonnet-4 as the default configuration for macro benchmarks.
@@
Ablation Study. We conduct an ablation to know how the adversarial design in ACToR drives correctness gains on the micro-benchmark. We analyze the improvement offered by two design choices: (a) having an adversarial discriminator in the workflow to guide the translator, and (b) forcing the discriminator agent to use the fuzzing script we provided. We compare three variants of ACToR: coverage-baseline, ACToR-NoFuzz, and the full ACToR (with fuzzing). The results summed across programs at the final iteration are reported in Figure 3. Entry $(\text{row},\text{column})$ shows the relative pass rate of the row method on the column method’s tests. All three methods largely pass their own tests as expected.
@@
Figure 3: The relative comparison among 3 different translation methods on Claude Code. Entry $(\text{row},\text{column})$ is the relative pass rate of the row’s translation on the column method’s tests.
@@
On coverage-baseline tests, ACToR-NoFuzz and the full ACToR reach 92% and 88% passing rates, respectively; in the reverse direction, coverage-baseline achieves less than 70% passing rate on tests from either ACToR variant. This indicates that adversarial discriminators produce tests that more effectively expose semantic mismatches and result in more correct translations. Between ACToR variants, the full ACToR fails more on its own tests, but attains 82% vs. 75% when cross-tested against ACToR-NoFuzz. This suggests that the discriminating agent uses the fuzzing script to surface more challenging bugs in the translated code. We also measure the final test coverage of each method on the C code. The result is shown at the bottom of Figure 3. The coverage baseline works as expected and achieves the highest line coverage on the C code than two other ACToR variants. However, it is worse in correctness. This further supports the conclusion that adversarial translation is more efficient in improving correctness compared with purely sampling new test cases based on coverage.
@@
On our micro benchmark, the full ACToR system substantially improves translation correctness compared with the naive baseline. The adversarial discriminator is crucial to improving correctness, and the fuzzing tool further strengthens the discovery of mismatches.
@@
Figure 4: The relative pass rate when cross-comparing ACToR and coverage baseline (Cov-Base) on Claude Code at iteration 10. For each program, the left bar shows evaluating the translation from ACToR on tests generated in coverage-baseline; the right bar is evaluating the translation from coverage-baseline on ACToR tests. The length of each program in LoC is presented next to the program name.
@@ 4.3 Evaluation on Macro Benchmark
We compare ACToR results with the coverage baseline to examine the improvement achieved in correctness, using the relative pass rate metric. We use the Claude Code agent and Claude-Sonnet-4 for this experiment. In considering the necessary system calls used in the translated code, we discourage the use of unsafe by instructing in prompts, but do not enforce full safe Rust in the middle of the translation loop. We enforce 100% safe Rust at the end of the translation in ACToR, as an additional simple post-processing agent call.
The cross-comparison result at the $10^{th}$ iteration [^4] On 3 programs ‘sync‘, ‘users‘, and ‘mknod‘, both translation methods abort before 10 iterations after multiple reruns. We therefore compare them on the maximum available iteration. is shown in Figure 4. The left bar for each program is the relative pass rate of testing ACToR’s translation on the test cases from the coverage baseline. The right bar shows the reverse: relative pass rate of the coverage baseline against ACToR’s test cases. We see that ACToR outperforms the coverage baseline on 55/57 programs on the relative pass rate. ACToR gets 100% relative pass rate on 26 programs when running on the tests from the coverage baseline, while the coverage baseline can only get full pass on 1 program. On 57 programs, ACToR achieves an average 93.9% relative pass rate, 18.9% higher than the 75.0% relative pass rate obtained by the coverage baseline. The cross-comparison confirms that adversarial agentic design in ACToR plays a significant role in improving the translation correctness. In an absolute (non-relative) sense of correctness, by taking the union of the test cases from both ACToR and the coverage baseline, we get a test set of an average 90.6% coverage. ACToR achieves a 95.3% pass rate on it. This demonstrates the utility of ACToR on translating real-world projects with high correctness. We further analyze the only benchmark program, specifically the pr program, for which ACToR has lower correctness than the coverage baseline in the Appendix A.
ACToR automatically translates the BSDCoreUtil benchmark with high correctness, outperforming the coverage baseline on 5555/$57$ programs and achieving an average of $93.9\%$ relative pass rate.
@@ 5 Conclusion
We present ACToR, an agent-based automatic translation approach that aims for functionally correct C to Rust translations. Our method introduces a collaborative pipeline in which two agents iteratively refine the translations. At each iteration, the discriminant agents detect the semantic mismatches between the source code and the translated program, and report them as failing test cases. The translator agents then improve the translation to pass the current test set. Our experiments demonstrate the power of ACToR’s adversarial agentic setup to produce translations with high correctness, surpassing the quality of translation as compared to non-adversarial setups. A noteworthy advantage that ACToR brings to the translation landscape is the potential for large-scale translation with minimal or no human intervention, making ACToR the first approach of this kind.
@@
The artifact containing detailed instructions, benchmark programs, and evaluation scripts will be released soon. Section 3 describes the algorithmic design of ACToR, including assumptions and iteration parameters, while Section 4 outlines our evaluation methodology and benchmark datasets. Appendix B includes the prompt template used for the translator and discriminator agents. Together, these resources enable independent verification of both our experimental setup and results.
Appendix A Case Study
ACToR outperformed coverage baseline on 55/57 programs in the macro benchmark, while failing on one program name ‘pr‘. We further analyze the reason behind this failure. The correctness gap is caused by the corner difference between the environment where we run ACToR and the validation environment when we test the translation, due to an imperfect testing harness. The translator agent claims that all tests were passed during the translation procedure, but actually got timeouts on most cases during validation. After manually adjusting the code (1 line edit) to adapt to the validation environment, the translation achieves an $84\%$ relative pass rate, higher than the coverage baseline.
(ACToR/ Coverage Baseline) Translator Agent Prompt Task Description: ‘‘‘ Please translate the C program in the current folder to safe Rust (no unsafe code). Please create a xxx_rust folder in the current folder, where xxx is the name of the C program (you can see from the Makefile). Please use Cargo to build the Rust program. Once you have the compiled Rust program (typically in target/debug/xxx‘), please run testcmp.sh to compare the output with the original C program. Please run testcmp.sh --help to understand how to use the test script. Please debug to make sure the translation passes all the tests. ‘‘‘ Please start the translation now.
Figure 5: The task prompt for the translator agent.
(Coverage Baseline) Test Generator Agent Prompt — Adding Tests Task Description: ‘‘‘ Please add more tests for the C program in the current folder. The goal is to increase the line coverage while keeping tests valid per our definition. GOALS OF ADDING TESTS: 1. Increase the line coverage. 2. Avoid duplication with existing tests; focus on new, meaningful scenarios and edge cases. 3. Ensure validity: Run ‘./testcmp.sh compare ./xxx and confirm the C program passes all tests. Internally, this means two compiled versions of the C program (‘./xxx.ref and ‘./xxx‘) must produce identical results, ensuring issues are fixable (not nondeterminism). RESOURCES FOR ADDING TESTS: - C source code files: ‘*.c and ‘*.h for understanding intended behavior. - Makefile for building (‘make all‘) and cleaning (‘make clean‘). - testsXX.jsonl‘: existing tests and format examples. - testcmp.sh for running comparisons: ‘./testcmp.sh compare <BINARY>‘. To get the coverage of tests, run ‘./testcmp.sh coverage‘. - $x$.gcov file after collecting coverage, to check which lines are not covered. TASK: Add 3 new test cases to testsXX.jsonl where $X$ is the largest index among all testsXX.jsonl files. <...Test formatting requirements> <...Detailed documentation for the input format of the test runner> ‘‘‘ Now please add 3 more tests as requested above.
Figure 6: The task prompt for the test generator agent of the coversge baseline.
(ACToR) Discriminator Agent Prompt — Adding Tests Task Description: ‘‘‘ Please add more tests for the C program in the current folder. The goal is to reveal behavioral differences between the current Rust translation and the C reference while keeping tests valid per our definition. GOALS OF ADDING TESTS: 1. Expose divergences between Rust and C behavior (Rust should fail where C passes). 2. Avoid duplication with existing tests; focus on new, meaningful scenarios and edge cases. 3. Ensure validity: Run ‘./testcmp.sh compare ./xxx and confirm the C program passes all tests. Internally, this means two compiled versions of the C program (‘./xxx.ref and ‘./xxx‘) must produce identical results, ensuring issues are fixable (not nondeterminism). RESOURCES FOR ADDING TESTS: - C source code files: ‘*.c and ‘*.h for understanding intended behavior. - Makefile for building (‘make all‘) and cleaning (‘make clean‘). - testsXX.jsonl‘: existing tests and format examples. - testcmp.sh for running comparisons: ‘./testcmp.sh compare <BINARY>‘. To get the coverage of tests, run ‘./testcmp.sh coverage‘. IF {\$allow_fuzz} - fuzzer_template.py for fuzzing: python fuzzer_template.py <BINARY1> <BINARY2>‘. You can read its code to understand how to generate tests. ENDIF - Rust translation project if present (e.g., ‘<prog>_rust with Cargo). Build with cargo build and try ‘./testcmp.sh compare <path_to_rust_binary>‘ to probe differences. - $x$.gcov file after collecting coverage, to check which lines are not covered. TASK: Find 3 new input cases that expose mismatches between C code and Rust translation. Add these 3 new test cases to testsXX.jsonl where $X$ is the largest index among all testsXX.jsonl files. <...Test formatting requirements> IF {\$allow_fuzz} <...Details on the usage of the fuzzer script> ENDIF <...Detailed documentation for the input format of the test runner> ‘‘‘ Now please add 3 more tests as requested above.
Figure 7: The task prompt for the discriminator agent of ACToR. The usage of the fuzzing tool is controlled by ‘\$allow_fuzz‘ in the prompt.
Appendix B Prompt Template
@@
For Mini-SWE-Agent, we keep the system prompt that describes the agent workflow unchanged as the original one during all the experiments. For Claude Code, we begin directly with the task described in the user message, without any additional text preceding it. The task prompt structure for the translator agent is shown in Figure 5. This prompt is shared among all translation methods during evaluation. The prompt structure for the test generation agent of the coverage baseline is shown in Figure 6. The prompt for the discriminator agent for ACToR is shown in Figure 7. The fuzzing mode is controlled by the ‘allow_fuzz‘ section in the prompt.

### v2

@@ 4.2 Evaluation Result on Micro Benchmark
Figure 2: Overall correctness (pass rate) achieved by ACToR on micro benchmark across different settings compared with naive baseline, on 5 different agent-model settings.
@@
Correctness & Adaptability. We first test ACToR on the micro benchmark with different agents and LLMs. To analyze the correctness improvement, we compare ACToR approach with a naive baseline on pass rate. The experiment results per program are shown in Figure 2. All translations from ACToR and the naive baseline are 100% safe code and free of unsafe blocks. ACToR consistently improves the translation correctness on all 5 agent-model settings. With Claude Code and Claude-Sonnet-4.5, the validation pass rate rises from an average of 89.2% (one-shot naive baseline) to 98.2% after 10 iterations. For Mini-SWE-Agent, pass rates increase from 92.5% to 97.2% with Claude-Sonnet-4.5 and from 84.1% to 86.8% with GPT-5 mini on average. Across agent-LLM settings, the closed-source Claude Code, together with the Claude-Sonnet-4.5, achieves the highest final pass rate. We therefore use Claude Code + Claude-Sonnet-4.5 as the default configuration for the following experiments. When using Claude Code and Claude-Sonnet-4.5, it takes 411 million tokens to finish 6 programs, which costs \$201 USD in total. The average time cost for ACToR to finish one program is approximately 2.45 hours [^4] The token consumption is computed by summing up the input, output, and cached tokens reported in the API response messages. When computing the time cost, we first compute the average time cost per iteration, then times it by 10 as an approximation of the total runtime for translating one program..
@@
Ablation Study on Design. We conduct an ablation to know how the adversarial design in ACToR drives correctness gains on the micro-benchmark. We analyze the improvement offered by two design choices: (a) having an adversarial discriminator in the workflow to guide the translator, and (b) forcing the discriminator agent to use the fuzzing script we provided.
@@
Figure 3: The relative comparison among 3 translation methods on Claude Code with Claude-Sonnet-4.5. Entry $(\text{row},\text{column})$ is the relative pass rate of the row’s translation on the column method’s tests.
@@
We compare three variants of ACToR: coverage-baseline, ACToR-NoFuzz, and the full ACToR (with fuzzing). The results summed across programs at the final iteration are reported in Figure 3. Entry $(\text{row},\text{column})$ shows the relative pass rate of the row method on the column method’s tests. All three methods largely pass their own tests as expected. On coverage-baseline tests, ACToR-NoFuzz and the full ACToR reach 93% and 96% passing rates, respectively; in the reverse direction, coverage-baseline achieves less than 60% passing rate on tests from either ACToR variant. This indicates that adversarial discriminators produce tests that more effectively expose semantic mismatches and result in more correct translations. Between ACToR variants, the full ACToR attains 82% vs. 81% when cross-tested against ACToR-NoFuzz. This suggests that the discriminating agent uses the fuzzing script to surface more challenging bugs in the translated code. We also measure the final test coverage of each method on the C code. The result is shown at the bottom of Figure 3. The coverage baseline works as expected and achieves the highest line coverage on the C code than two other ACToR variants. However, it is worse in correctness. This further supports the conclusion that adversarial translation is more efficient in improving correctness compared with purely sampling new test cases based on coverage. Finally, we compare the financial costs of completing all six programs: the coverage baseline costs \$95 USD, while ACToR-NoFuzz and the full ACToR cost \$240 USD and \$211 USD, respectively. To better understand cost-effectiveness, we ran an equal-cost experiment in which we extended the coverage baseline to 25 iterations, matching the approximate cost of running ACToR for 10 iterations. Even under equal budget, the coverage baseline does not close the performance gap: its relative pass rate (165/267 = 61.8%) remains far below that of ACToR (515/537 = 95.9%). This demonstrates that ACToR delivers substantially higher correctness per dollar spent. Additional details are shown in the Appendix B.
Figure 4: The validation pass rate of ACToR on different configurations. Pass rate versus the number of new test cases added per iteration (Left). Pass rate versus the number of iterations (Right).
[…]
Stability of Results. To assess the stability of ACToR ’s translation results, we run ACToR three times using the default settings (i.e., 15 initial tests, 10 iterations, and 3 tests per iteration). We record the pass rate of each trial on the validation set. For each program, we compute the sample standard deviation of the three pass rates and then average these standard deviations across all programs. On average, the three runs achieve a pass rate of 97.0%, with a deviation of 1.9 percentage points.
@@
On the micro-benchmark, our system substantially and stably improves translation correctness compared with the naive baseline. The adversarial discriminator is crucial to improving correctness, and the fuzzing tool further strengthens the discovery of mismatches.
@@
Figure 5: The relative pass rate when cross-comparing ACToR and coverage baseline (Cov-Base) on Claude Code with Claude-Sonnet-4.5 at iteration 10. For each program, the left bar shows evaluating the translation from ACToR on tests generated in coverage-baseline; the right bar is evaluating the translation from coverage-baseline on ACToR’s tests. The length of each program in LoC is presented next to the program name.
@@ 4.3 Evaluation on Macro Benchmark
We compare ACToR results with the coverage baseline to examine the improvement achieved in correctness, using the relative pass rate metric. We use the Claude Code agent and Claude-Sonnet-4.5 for this experiment [^5] We run macro evaluation on both Claude-Sonnet-4 and Claude-Sonnet-4.5. Due to space limitations, we put the result of Claude-Sonnet-4 in Appendix C for reference. The results are consistent between two trials.. All the final translations output by ACToR are free of unsafe blocks.
The cross-comparison result at the $10^{th}$ iteration is shown in Figure 10. The left bar for each program is the relative pass rate of testing ACToR’s translation on the test cases from the coverage baseline. The right bar shows the reverse: relative pass rate of the coverage baseline against ACToR’s test cases. We see that ACToR outperforms the coverage baseline on 54/57 programs on the relative pass rate. ACToR gets 100% relative pass rate on 32 programs when running on the tests from the coverage baseline, while the coverage baseline can only get full pass on 3 programs. On 57 programs, ACToR achieves an average 96.4% relative pass rate, 25.1% higher than the 71.3% relative pass rate obtained by the coverage baseline. The cross-comparison confirms that adversarial agentic design in ACToR plays a significant role in improving the translation correctness. In an absolute (non-relative) sense of correctness, by taking the union of the test cases from both ACToR and the coverage baseline, we get a test set of an average 90.1% coverage. ACToR achieves a 95.9% pass rate on it. This demonstrates the utility of ACToR on translating real-world projects with high correctness. The total financial costs to translate 57 programs using coverage-baseline and ACToR are \$808 USD and \$1634 USD, respectively. On average, ACToR consumes 57 million tokens per program, with more than 97% of them being cache-read tokens. The average time cost for ACToR to finish one program is about 2.5 hours.
ACToR automatically translates the BSDCoreUtil benchmark with high correctness with reasonable costs, outperforming the coverage baseline on 54/$57$ programs and achieving an average of 96.4% relative pass rate.
@@ 5 Conclusion
We present ACToR, an agent-based automatic translation approach that aims for functionally correct C to Rust translations. Our method introduces a collaborative pipeline in which two agents iteratively refine the translations. At each iteration, the discriminant agents detect the semantic mismatches between the source code and the translated program, and report them as failing test cases. The translator agents then improve the translation to pass the current test set. Our experiments demonstrate the power of ACToR’s adversarial agentic setup to produce translations with high correctness, surpassing the quality of translation as compared to non-adversarial setups. A noteworthy advantage that ACToR brings to the translation landscape is the potential for translating a large number of command-line programs with minimal or no human intervention, making ACToR the first approach of this kind.
@@
The artifact containing detailed instructions, benchmark programs, and evaluation scripts will be released soon. Section 3 describes the algorithmic design of ACToR, including assumptions and iteration parameters, while Section 4 outlines our evaluation methodology and benchmark datasets. Appendix A includes the prompt template used for the translator and discriminator agents. Together, these resources enable independent verification of both our experimental setup and results.
Appendix A Prompt Template
@@
For Mini-SWE-Agent, we keep the system prompt that describes the agent workflow unchanged as the original one during all the experiments. For Claude Code, we begin directly with the task described in the user message, without any additional text preceding it. The task prompt structure for the translator agent is shown in Figure 6. This prompt is shared among all translation methods during evaluation. The prompt structure for the test-generation agent used in the coverage baseline is shown in Figure 7. The prompt structure for the discriminator agent for ACToR is shown in Figure 8. The fuzzing mode is controlled by the ‘allow_fuzz‘ section in the prompt.
Appendix B Equal-Cost Comparison on Micro Benchmark
Running all six programs in the micro benchmark costs \$95 USD for the coverage baseline and \$211 USD for ACToR. To better understand cost-effectiveness, we ran an equal-cost experiment in which we extended the coverage baseline to 25 iterations to match the cost of running ACToR for 10 iterations. We cross-compare the resulting translations using relative pass rate. The two methods complete different numbers of iterations, resulting in different numbers of test cases. The total USD cost for the coverage baseline to finish 25 iterations is \$220 USD, roughly the same as ACToR to finish 10 iterations (\$211 USD). However, even under equal budget, the coverage baseline does not achieve comparable correctness. Its relative pass rate (165/267 = 61.8%) remains far below that of ACToR (515/537 = 95.9%). In absolute terms, ACToR fails only 22 out of 537 tests, while the coverage baseline fails 102 out of 267. This demonstrates that ACToR delivers substantially higher correctness per dollar spent. Additional details are shown in the appendix.
Appendix C Macro Experiments with Claude Code and Claude-Sonnet-4
We also run ACToR with Claude-Sonnet-4 to translate the macro-benchmark. In this trial, we use a slightly different implementation that discourages the use of unsafe, but does not enforce full safe Rust in the middle of the translation loop. We enforce full safe Rust at the end of the translation, as a post-processing step. This post-processing is performed via a single agent call that is required to eliminate all unsafe blocks while preserving the functionality of translated Rust code. The prompt for this post-processing step is shown in Figure 9. We emphasize that this strategy slightly differs from the approach used in the main paper when run ACToR with Claude-Sonnet-4.5, where we enforce full safety at every iteration of the translation loop. All other experimental settings in this trial are identical to those described in the main text. The results are presented in Figure 10. With Claude-Sonnet-4, ACToR outperforms the coverage baseline on 55/57 programs in terms of relative pass rate. ACToR achieves full pass on 26 programs using the tests from the coverage baseline. Across all 57 programs, ACToR attains an average relative pass rate of 93.9%, which is 18.9% higher than the 75.0% pass rate of the coverage baseline. ACToR achieves a 95.3% pass rate on the union of the test cases from both ACToR and the coverage baseline. This result aligns with the findings on Claude-Sonnet-4.5, demonstrating that ACToR’s performance is stable and generalizes across different model configurations.
Figure 10: The relative pass rate when cross-comparing ACToR and coverage baseline (Cov-Base) on Claude Code with Claude-Sonnet-4 at iteration 10. For each program, the left bar shows evaluating the translation from ACToR on tests generated in coverage-baseline; the right bar is evaluating the translation from coverage-baseline on ACToR’s tests. The length of each program in LoC is presented next to the program name.